OpenStudy (anonymous):

Paul’s Pool Service charges $340 to open a pool and charges $28 each week for pool maintenance. Nelson’s Pool Service charges $280 to open a pool and charges $32 each week for pool maintenance. The system that models this situation is given, where c is the cost of pool maintenance and w is the number of weeks. c = 340 + 28w c = 280 + 32w The solution to the system is (15, 760). Which interpretation correctly describes the solution to the system of equations?

OpenStudy (anonymous):

Paul’s Pool Service charges $340 to open a pool and charges $28 each week for pool maintenance. Nelson’s Pool Service charges $280 to open a pool and charges $32 each week for pool maintenance. The system that models this situation is given, where c is the cost of pool maintenance and w is the number of weeks. c = 340 + 28w c = 280 + 32w The solution to the system is (15, 760). Which interpretation correctly describes the solution to the system of equations? A. At 15 wk of pool cleanings, a customer will reach the maximum payment of $535. B. At 15 wk, the cost for pool maintenance is the same for both companies. Both companies will charge $760. C. Paul’s Pool Service charges more money in the fifteenth week by charging $760. D. Nelson’s Pool Service charges more money in the fifteenth week by charging $760.

OpenStudy (anonymous):

@logancoop

OpenStudy (anonymous):

Its B